OpenStudy (anonymous):

Can anyone help evaluate this for me? -8theta^2cos(theta/4) and theta is pi/2

OpenStudy (anonymous):

\[-8\theta ^{2}\cos (\theta \div4) \]theta = pi/2

OpenStudy (anonymous):

-18.236651